OpenStudy (anonymous):

Simplify sqrt24 - 3sqrt45 + 2sqrt20

OpenStudy (callisto):

\[\sqrt{24} - 3\sqrt{45} + 2\sqrt{20}\]\[=\sqrt{2^2\times6} - 3\sqrt{3^2\times 5} + 2\sqrt{2^2\times 5}\]\[=2\sqrt{6} - 9\sqrt{5} + 4\sqrt{5}\]Can you do it? Group the similar terms